Gino’s Restaurant is a popular restaurant in Boston, Massachusetts. The owner of the restaurant has been trying to better understand costs at the restaurant and has hired a student intern to conduct an activity-based costing study. The intern, in consultation with the owner, identified the following major activities:

  Activity Cost Pool Activity Measure
  Serving a party of diners      Number of parties served
  Serving a diner      Number of diners served
  Serving drinks      Number of drinks ordered

Some costs, such as the cost of cleaning the linens that cover the restaurant's tables, vary with the number of parties served. Other costs, such as washing plates and glasses, depends on the number of diners served or the number of drinks served.

  

  Data concerning these activities are shown below.

   

Serving a Party Serving a Dinner Serving Drinks Total
  Total cost $51,000 $124,000 $95,200 $270,200
  Total activity 6,000 parties 20,000 diners 56,000 drinks

   Prior to the activity-based costing study, the owner knew very little about the costs of the restaurant. She knew that the total cost for the month was $270,200 and that 20,000 diners had been served. Therefore, the average cost per diner was $13.51 ($270,200 ÷ 20,000 diners = $13.51 per diner).

Required:

1.

Compute the activity rates for each of the three activities. (Round your answers to 2 decimal places.)

     

2. According to the activity-based costing system, what is the total cost of serving each of the following parties of diners? (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

   

a. A party of four diners who order three drinks in total.
b. A party of two diners who do not order any drinks.
c. A party of one diner who orders two drinks.


3. Convert the total costs you computed in part (1) above to costs per diner. In other words, what is the average cost per diner for serving each of the following parties? (Do not round intermediate calculations. Round your final answers to 2 decimal places.)

   

a. A party of four diners who order three drinks in total.
b. A party of two diners who do not order any drinks.
c. A party of one diner who orders two drinks.
Answer #1

Ans:

Activity cost Pool Total
1) Total cost Activity Rate
Serving a party of diners 51,000 6,000 8.5 per party
Serving a diner 1,24,000 20,000 6.2 per diner
Serving drinks 95,200 56,000 1.7 per drink
2)
a) Activity cost Pool Total Total cost
Rate Activity
Serving a party of diners 8.5 1 8.5
Serving a diner 6.2 4 24.8
Serving drinks 1.7 3 5.1
Total cost 38.4
b)
Activity cost Total Total Cost
pool Rate Activity
Serving a party of diners 8.5 1 8.5
Serving a diner 6.2 2 12.4
Serving drinks 1.7 0 0
Total cost 20.9
c)
Activity cost Total Total cost
pool Rate Activity
Serving a party of diners 8.5 1 8.5
Serving a diner 6.2 1 6.2
Serving drinks 1.7 2 3.4
Total cost 18.1
3) a) Average cost of dinner=Total cost/No of dinners=38.4/4=9.6
b) Average cost of dinner=Total cost/No of dinners=20.9/2=10.45
c) Average cost of dinner=Total cost/No of dinners=18.1/1=18.1